OpenStudy (anonymous):

List the solutions between 0, and 2pi for sin(x/3)= sqrt(2)/2

OpenStudy (mertsj):

sin of pi/4 and 3pi/4 is sqrt(2)/2 So it follows that:

OpenStudy (mertsj):

\[\frac{x}{3}=\frac{\pi}{4}\] and \[x=\frac{3\pi}{4}\]

OpenStudy (mertsj):

Also \[\frac{x}{3}=\frac{3\pi}{4}\] so \[x=\frac{9\pi}{4}\]

OpenStudy (mertsj):

But 9pi/4 is not between 0 and 2pi so there is only 1 solution.